OpenStudy (anonymous):

1.The distance on the number line from -10 to -2 equals 5x – 12. What is the value of x? 2.The distance on the number line from 5 to 12 equals 3x – 2. What is the value of x?

OpenStudy (anonymous):

Is this question from a test ?

OpenStudy (anonymous):

why -10-12 ? shouldnt it be -10 +2 ?

OpenStudy (igreen):

Huh..? It says: "The distance on the number line from -10 to -2" So it's not + 2, it's - 2.

OpenStudy (anonymous):

Yes but isnt -10-(-2) =-10 +2 ?

OpenStudy (kittiwitti1):

\[| -10 - (-2) | = 5x - 12\]\[| -10 + 2 | = 5x -12~which~becomes~ |-8| = 5x-12~and~then~8=5x-12.\]Solve for x.

OpenStudy (anonymous):

-10 -2=5x-12 -12 + 12 = 5x 0= 5x 0= x ??

OpenStudy (kittiwitti1):

... no, did you read my explanation?

OpenStudy (anonymous):

Yes sorry I think I understand ! Thank you !

OpenStudy (kittiwitti1):

Okay. xD You still need help on the second one or no? c:

OpenStudy (anonymous):

8=5x−12 20 = 5x 4= x

OpenStudy (kittiwitti1):

Wut. xD

OpenStudy (kittiwitti1):

Oh! Yeah, you got it right! Good job ^^

OpenStudy (anonymous):

I will try to do it alone thank you

OpenStudy (anonymous):

@kittiwitti1 |5−12|=3x−2 |-7|=3x−2 7=3x-2 9=3x 3= x Is that correct ?